The demand function for good X is 𝑄π‘₯ 𝑑 = π‘Ž βˆ’ 𝑏𝑃π‘₯ + 𝑐𝑀 + 𝑒 . Where 𝑃π‘₯ is the price of good X and M is income . Least squares regression reveals that Γ’ = 8.27, bΛ† = 2.14, cΛ† = 0.36, 𝜎Ò = 5.32, πœŽπ‘Λ† = 0.41, and πœŽπ‘Λ† = 0.22. The R-squared is 0.35. a. Compute the t-statistic for each of the estimated coefficients. (4mks) b. Determine which (if any) of the estimated coefficients are statistically different from zero. (4mks) c. Explain, in plain words, what the R-square in this regression indicates. (


1
Expert's answer
2022-02-09T11:04:52-0500

"Qxd=a-bPx+cm+e"

Fitting in the estimated coefficients

"Qxd=8.27-2.14Px+0.36M+e"

"ta=\frac{a}{da}=\frac{8.27}{5.32}=1.55"

"tb=\frac{b}{db}=\frac{2.14}{0.41}=5.22"

"tc=\frac{c}{dc}=\frac{0.36}{0.22}=1.64"

Where ta, tb and tc represents the t test values for a, b and c respectively.

- Only the estimated coefficients of b is statistically different from zero. This is because it's t-value (5.22) is greater than the table t-value= 2.776.

- The R-square 0.35 explains 35% of the total valuation of demand for good X in relation to price and income. The remaining 65% is attributed to the error term and other factors affecting demand.